SEP SPECTRA OF SCCl_{2}

\underline{SANDRA LEE} and MARTIN GRUEBELE, Department of Chemistry, University of Illinois at Urbana-Champaign, Urbana IL 61801.

Stimulated emission pumping spectra with 0.3 cm^{−1} resolution provide a means of studying intramolecular vibrational energy redistribution (IVR). The Ã(^{1}A_{2}) → X̃(^{1}A_{1}) transition of SCCl_{2} from several different origins maps out the 7000-8000 cm^{−1} region in the ground state. An analysis of approximately 200 features in this region exhibits clumping of spectral features which can provide information about the fine structure of IVR .